# Analytical Inverse Kinematic Solution for ”Moz1” NonSRS 7-DOF Robot arm with novel arm angle

### v1

@@ I Introduction
Research on light-weight redundant manipulators, has grown in various directions like human robot interaction [kirschner2021involuntary] or machine learning [petrik2021learning]. While some of these applications use the inverse kinematics solver which is integrated in the robot’s control system, others use a numerical solver. The former solution is only available in position control mode and the code of the robot control unit is closed source and information on the used algorithms and therefore the redundancy resolution can hardly be found. Numerical solvers, on the other hand, always come with a risk of violating real-time constraints or not finding a solution at all, although solutions exist in the workspace[beeson2015trac].
Finding an analytical solution for the IKP of serial robotic manipulators is one of the fundamental basic problems in robotics and there is still no general solution available. There exist many analytical solutions for specific classes of robots like 7R anthropomorphic robot arms with spherical shoulder, spherical wrist(SRS structure), and three intersecting joint axes in the elbow [liu2017analytical], [bongardt2017inverse]. Although NonSRS structure manipulation(like Franka Emika Panda) is also a 7R anthropomorphic robot arm, the previously mentioned solutions do not apply for the robot since there are linear offsets in its elbow and wrist joints. These offsets make the analytical solution more complex compared to other anthropomorphic arms like the KUKA iiwa, where all joints are aligned in the zero configuration. More details on the robot’s structure are given in section II.
For SRS strucure, The IKP is solved based on the SEW angle as an redundancy represent[shimizu2008analytical], The IKP of some non-zero link offsets robot(NonSRS sturcture) could also be analytically solved based on the SEW angle method[an2014analytical], it required that the shoulder and wrist must be static during self-motion(we call it static-SW 7dof robot) However, for those nonstatic-SW robot, To the best of the author’s knowledge, there is no analytical IKP implementation available. This paper provides both the analytical solution as well as a real-time safe implementation for the nonstatic-SW 7dof robot. More details about inverse kinematic are given in section III, first in a specific configuration then promote it to general case.
To validate the solutions presented in this paper, a C++ version of the approach has been implemented . Details on the implementation and used libraries with visualization of results are presented in section IV. A conclusion of this work is given in section V, where also some other robots with open IKP are mentioned and possible future work based on this solution is illustrated.
@@ II KINEMATIC STRUCTURE
The Moz1 robot consists of eight links connected by seven joints, which make the serial robot a redundant manipulator. figure 1 show the MDH parameters and figure 2 illustrates its kinematic structure. The shoulder is constructed similar to other anthropomorphic arms like the KUKA iiwa, i.e., it is also a spherical shoulder. the first six joint is Similar to KUKA iiwa while there is an offset bettwen sixth joint and seventh joint, so there axis are not intersecting. which can be seen in figure 2. the offset awr = 0.0905 m .
@@ III-B Special configuration
The ik method above is only valid for a very small subset of all possible end effector poses in the work space where the wrist center C is located on the base z-axis and the 7th joint axis z7 is located in the base xz-plane. However, a closer look on the robot’s kinematics shows that all possible end effector poses can be reduced to one of those special poses. [tittel2023analysis] shows an example of how general poses can be reduced to a special pose which is similar to our case. By freeing the first 3 joint (sphere joint). it would rotate the robot around the sphere until it become a special configuration, it is easy to see that q4 q5 q6 q7 is state in such motion. so we could get the q4 to q7 with the same method, then using q4 to q7 to solve a sphere rotation to get q1 q2 q3. the key step is get the 3 dof $d_{sc}$, q and al from given endpose;
@@ IV Singularity analysis
the singularity analysis is also a basic part in robot, especilly in redundency robot, there are no simple way to get all singularities.Here we use the reciprocity-based method[nokleby2001reciprocity] and give the result of the singularity:

### v2

@@ I Introduction
Research on light-weight redundant manipulators, has grown in various directions like human robot interaction [4] or machine learning [7]. While some of these applications use the inverse kinematics solver which is integrated in the robot’s control system, others use a numerical solver. The former solution is only available in position control mode and the code of the robot control unit is closed source and information on the used algorithms and therefore the redundancy resolution can hardly be found. Numerical solvers, on the other hand, always come with a risk of violating real-time constraints or not finding a solution at all, although solutions exist in the workspace[2].
Finding an analytical solution for the IKP of serial robotic manipulators is one of the fundamental basic problems in robotics and there is still no general solution available. There exist many analytical solutions for specific classes of robots like 7R anthropomorphic robot arms with spherical shoulder, spherical wrist(SRS structure), and three intersecting joint axes in the elbow [5], [3]. Although NonSRS structure manipulation(like Franka Emika Panda) is also a 7R anthropomorphic robot arm, the previously mentioned solutions do not apply for the robot since there are linear offsets in its elbow and wrist joints. These offsets make the analytical solution more complex compared to other anthropomorphic arms like the KUKA iiwa, where all joints are aligned in the zero configuration. More details on the robot’s structure are given in section II.
In case of the 7dof robot, for SRS strucure, the IKP is solved based on the SEW angle as an redundancy represent[8], The IKP of some non-zero link offsets robot(NonSRS sturcture) could also be analytically solved based on the SEW angle method[1], it required that the shoulder and wrist must be static during self-motion(we call it static-SW 7dof robot) However, for those nonstatic-SW robot, To the best of the author’s knowledge, there is no analytical IKP implementation available.
This paper provides both the analytical solution as well as a real-time safe implementation for the nonstatic-SW 7dof robot. More details about inverse kinematic are given in section III, first in a specific configuration then promote it to general case.
In addition, the singularity of this robot is analyzed and classified with reciprocity-based method, details on the singularity are presented in section IV. A conclusion of this work is given in section V, where also some other robots with open IKP are mentioned and possible future work based on this solution is illustrated.
@@ II KINEMATIC STRUCTURE
The Moz1 robot consists of eight links connected by seven joints, which make the serial robot a redundant manipulator. figure 1 show the MDH parameters and figure 2 illustrates its kinematic structure. The shoulder is constructed similar to other anthropomorphic arms like the KUKA iiwa, i.e., it is also a spherical shoulder. the first six joint is Similar to KUKA iiwa while there is an offset bettwen sixth joint and seventh joint, so their axis are not intersecting. which can be seen in figure 2. the offset awr = 0.0905 m .
@@ III-B Special configuration
The ik method above is only valid for a very small subset of all possible end effector poses in the work space where the wrist center C is located on the base z-axis and the 7th joint axis z7 is located in the base xz-plane. However, a closer look on the robot’s kinematics shows that all possible end effector poses can be reduced to one of those special poses. [9] shows an example of how general poses can be reduced to a special pose which is similar to our case. By freeing the first 3 joint (sphere joint). it would rotate the robot around the sphere until it become a special configuration, it is easy to see that q4 q5 q6 q7 is state in such motion. so we could get the q4 to q7 with the same method, then using q4 to q7 to solve a sphere rotation to get q1 q2 q3. the key step is get the 3 dof $d_{sc}$, q and al from given endpose;
@@ IV Singularity analysis
the singularity analysis is also a basic part in robot, especilly in redundency robot, there are no simple way to get all singularities.Here we use the reciprocity-based method[6] and give the result of the singularity:
@@ V IMPLEMENTATION AND RESULTS
Furthermore, we believe that the methods presented in the paper can inspire analytical inverse solutions with arm angle for Franka robots.